OpenStudy (anonymous):

help trying to sove this ... write an equation of the line containig the given point and parrellel to the given line express your answer in form of y=mx+b (8,9) x+2y=9

OpenStudy (anonymous):

y=x/2+5

OpenStudy (mathmate):

For the original line x+2y=9 to pass through (8,9), substitute x and y into the equation and replace the 9 by 8+2*9=26. The new line is therefore x+2y=26. Convert to the y=mx+b form: y=-x/2+13